Calculate the entropy change of 1.0 mol of an ideal gas that undergoes an isothermal transformation form an initial state of pressure pi = 1.50atm and volume 500cm^3 to a final state of pressure of = 0.09atm.
Answer #1

From Ideal Gas equation, PV=nRT
P=Pressure

V=Volume
n=Number of moles

T=Temperature

R=Universal Gas Constant=8.314 J/Kmol
So for an ideal gas undergoing isothermal transformation, Temperature is constant. So PV=a constant
or P1V1=P2V2

so 1.5*500/0.09=V2

V2=8333cm3
\Delta s=nRln\frac{V_{2}}{V_{1}}

By Substituting,\Delta s=1*8.314*ln2.81=23 J/K